Four years ago man was six times as old as his son but in five years time he will be only three times as old as his son, what are their present ages
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement, simplify
Four years ago man was six times as old as his son
m - 4 = 6(s - 4)
m - 4 = 6s - 24
m - 6s = -24 + 4
m - 6s = -20
but in five years time he will be only three times as old as his son,
m + 5 = 3(s + 5)
m + 5 = 3s + 15
m - 3s = 15 - 5
m - 3s = 10
Use elimination on these two equations
m - 3s = 10
m - 6s = -20
----------------subtraction eliminates m, find s
0 + 3s = 30
s = 30/3
s = 10 yrs is the son's present age
and
m - 3(10) = 10
m = 10 + 30
m = 40 yrs is Dad's age
:
:
Check this in the statement:
"Four years ago man was six times as old as his son"
40 - 4 = 6(10-4)
36 = 6(6)